Please help me with this one. ** An empty swimming pool can be filled in 10 hours. When full, the pool can be drained in 16 hours. How long will it take to fill the pool if the drain is left open?
Found 2 solutions by venugopalramana, checkley71:
Answer by venugopalramana(3286)

Please help me with this one. ** An empty swimming pool can be filled in 10 hours.
SO IN 1 HOUR WE CAN FILL 1/10 POOL
When full, the pool can be drained in 16 hours.
SO IN 1 HOUR 1/16 POOL WILL BE DRAINED
How long will it take to fill the pool if the drain is left open?
SO IN 1 HOUR 1/10 - 1/16 = (8-5)/80 = 3/80 POOL WILL BE FILLED
SO IT TAKES 80/3 = 26.67 HRS TO FILL THE POOL

Answer by checkley71(8403)

I'LL GIVE IT MY BEST SHOT!!!!!!!!!!!!!!!!!!!
ASSUMING THE POOL IS 5400 CUBIC FEET OF WATER THEN THE INPUT RATE IS 4800/10
480 GALLONS PER HOUR THE OUTLET IS 4800/16 OR 300 GAL PER HOUR THEREFORE THE
DIFFERENCE IN RATE IS 180 GALLONS PER HOUR THUS TO FILL THIS POOL WITH
4800 GALLONS IT WILL TAKE 4800/180 OR 26 2/3 HOURS